The length of a rectangle is 2 less than 3 times its width. If the area measures 65 square meters find the dimensions of the rectangle.
Answer by cleomenius(959)

A = L*W
(3x-2)(x) = 65
3x^2 -2x -65 = 0
(3x + 13)( x - 5)
x = -13/3, x = 5
3(5) - 2 = 13 M = length
5 m = Width.
Cleomenius.